# Automatic Forward Model Parameterization with Bayesian Inference of Conformational Populations

### v2

@@ Abstract
To quantify how well theoretical predictions of structural ensembles agree with experimental measurements, we depend on the accuracy of forward models. These models are computational frameworks that generate observable quantities from molecular configurations based on empirical relationships linking specific molecular properties to experimental measurements. Bayesian Inference of Conformational Populations (BICePs) is a reweighting algorithm that reconciles simulated ensembles with ensemble-averaged experimental observations, even when such observations are sparse and/or noisy. This is achieved by sampling the posterior distribution of conformational populations under experimental restraints as well as sampling the posterior distribution of uncertainties due to random and systematic error. In this study, we enhance the algorithm for the refinement of empirical forward model (FM) parameters. We introduce and evaluate two novel methods for optimizing FM parameters. The first method treats FM parameters as nuisance parameters, integrating over them in the full posterior distribution. The second method employs variational minimization of a quantity called the BICePs score that reports the free energy of “turning on” the experimental restraints. This technique, coupled with improved likelihood functions for handling experimental outliers, facilitates force field validation and optimization, as illustrated in recent studies (Raddi et al. 2023, 2024). Using this approach, we refine parameters that modulate the Karplus relation, crucial for accurate predictions of $J$-coupling constants based on dihedral angles ($\phi$) between interacting nuclei. We validate this approach first with a toy model system, and then for human ubiquitin, predicting six sets of Karplus parameters for ${{}^{3}\!J}_{H^{N}H^{\alpha}}$, ${{}^{3}\!J}_{H^{\alpha}C^{\prime}}$, ${{}^{3}\!J}_{H^{N}C^{\beta}}$, ${{}^{3}\!J}_{H^{N}C^{\prime}}$, ${{}^{3}\!J}_{C^{\prime}C^{\beta}}$, ${{}^{3}\!J}_{C^{\prime}C^{\prime}}$. Finally, we demonstrate that our framework naturally generalizes optimization to any differentiable forward model, such as those constructed with neural networks. This approach, which does not rely on any predetermined parameterization, enhances predictive accuracy and can be used for many applications.
@@ I Introduction
Model validation and refinement of structural ensembles against NMR observables critically depends on reliable forward models (FMs) that have been robustly parameterized to minimize FM error. In refining these FMs, it is important to consider random and systematic errors inherent to the experimental data, which in turn contribute to overall uncertainty of integrative models. It is usually challenging to obtain complete knowledge of such errors; instead, such errors typically need to be inferred from the data.
@@
To effectively refine FM parameters, we use BICePs to sample over the entire posterior distribution of FM parameters, enabling simultaneous ensemble reweighting and parameter refinement. Additionally, we show that by variational minimization of the BICePs score, we obtain the same result and show that the two approaches are equivalent, with each method requiring particular considerations. We first demonstrate our method’s effectiveness using a toy model and then apply it to optimize six distinct sets of Karplus parameters for the human protein ubiquitin. As a proof-of-concept for more complex models, we also train a neural network to learn $J$-coupling constants from dihedral angles using the BICePs score as a loss function. We then discuss how this idea could be applied to neural networks for chemical shift prediction.Han et al. (2024); Yang, Chakraborty, and White (2021); Meiler (2003); Shen and Bax (2010) Of course, our framework naturally extends to other empirical forward models with tunable parameters—for example, hydrogen-deuterium exchange (HDX) models with adjustable hydrogen-bonding and burial weights, or paramagentic relaxation enhancement (PRE) models with system-specific correlation times and scaling factors. These examples illustrate the generality and flexibility of our approach for refining a wide range of forward models to improve agreement between simulations and experiment.
@@ II Theory
In this section, we develop two different theoretical approaches to optimize forward models. In the first method, we include forward model parameters in the posterior, and sample the full posterior distribution of all parameters. In the second approach, we optimize the forward model by minimizing the BICePs score, a free energy-like quantity that can be used for variational optimization.
@@ II.1 Forward model optimization by posterior sampling of forward model parameters
The Good-Bad model accounts for systematic error due to outlier measurements.
Recent work has developed sophisticated likelihood models for BICePs, designed to detect outliers, that are robust in the presence of systematic errorRaddi et al. (2025). Raddi et al. (2025) developed two likelihood models for this purpose: the Student’s model and the Good-Bad model. In related work using BICePs to perform automated force field optimization, we used the Student’s model.Raddi and Voelz (2024) Here, in this work, we use the Good-Bad model, primarily because the derivatives of the Good-Bad model are less complicated than the Student’s model.
The Good-Bad likelihood modelRaddi et al. (2025) assumes that the level of noise is mostly uniform, except for a few erratic measurements. This limits the number of uncertainty parameters that need to be sampled, while still capturing outliers. Consider a model in which the uncertainties $\sigma_{j}$ for particular observables $j$ are distributed with respect to some typical uncertainty $\sigma^{B}$ according to a conditional probability $p(\sigma_{j}|\sigma^{B})$. We derive a posterior for the parameters of the $k^{th}$ FM having a single uncertainty parameter $\sigma^{B}$ by marginalizing over all $\sigma_{j}$,
[…]
@@
where $10\leq\omega<$ 1 describes the fraction of "bad" observables. Since the value of $\omega$ is unknown, it is treated as a nuisance parameter, and marginalized over its range. The resulting posterior is
[…]
@@
$\begin{split}u=\sum^{N}_{r=1}-\log(p(X_{r}))-N\sum^{N_{d}}_{j=1}\log\left[\frac{1-H(\sigma^{SEM}-\sigma_{0})}{2}\mathcal{N}\left(g_{j}(\mathbf{X},\theta^{(k)})\,\middle|\,d_{j},\sigma_{0}^{2}\right)+\frac{1-H(\sigma^{SEM}-\varphi\sigma_{0})}{2\varphi}\mathcal{N}\left(g_{j}(\mathbf{X},\theta^{(k)})\,\middle|\,d_{j},\varphi^{2}\sigma_{0}^{2}\right)\right].\end{split}$ (10)
@@ II.2 Forward model optimization by variational minimization of the BICePs score
As we show in Results and Discussion, treating the forward model parameters as nuisance parameters and sampling them as part of the full posterior is an efficient strategy to refine FM parameters. However, in the limit of large numbers of FM parameters, the dimensionality of the posterior may become unwieldy and may suffer from the “curse of dimensionality”.
@@
[…]
@@
It should be noted that in other applications of BICePs,Raddi et al. (2025); Raddi and Voelz (2024) the reference state for the BICePs score is defined using the $\lambda=0$ state for a series of a priors $p_{\lambda}(X)\sim[p(X)]^{\lambda}$, and the BICePs score is computed as the free energy of ($\lambda=0\rightarrow 1$) and ($\xi=0\rightarrow 1$) transformations. Here, since we are only interested in evaluating and/or parameterizing the likelihood functions, we set $p(X)$ to be uniform.
@@
For this purpose, gradient-based searches are very helpful. Conveniently, the derivative of the BICePs score $f(\theta)$ with respect to the FM parameters reduces to the Boltzmann-averaged values of the gradient of the $\partial u/\partial\theta$ :
[…]
@@ III Results
To evaluate the robustness of our algorithm in the presence experimental inaccuracies, we introduced random and systematic errors of varying magnitudes ($\sigma_{\text{data}}$) into the synthetic experimental scalar couplings. The performance of our Good-Bad likelihood model, a Gaussian likelihood model, and singular value decomposition (SVD) was compared under these conditions.
@@
[…]
@@
where $p(X)$ represents the true populations for state $X$, and $\phi_{0}$ is the phase shift of $-60^{\circ}$.
@@
Figure 2: Comparative analysis in performance of the Good-Bad likelihood model (red), a Gaussian likelihood model (blue), and singular value decomposition (SVD) using the "true" $\phi$ angles with synthetic experimental data. Here, we induced random and systematic error of varying magnitude ($\sigma_{\text{data}}$) to the experimental scalar couplings. Model performance was measured by computing RMSE (Hz) between the "true" scalar couplings and the couplings generated from the Karplus relations with predicted Karplus coefficients over 1,500 random perturbations to the experimental data, and represent the average of 100 BICePs calculations. Error bars represent the standard deviation. Predictions from SVD and the Gaussian likelihood model become notably less dependable when data incorporates errors, especially when $\sigma_{\text{data}}$ exceeds 0.5 Hz.
@@
As a method for forward model optimization, variational minimization of the BICePs score has advantages and disadvantages. This method is particularly advantageous for handling many FM parameters, offering a potential solution to the curse of dimensionality faced by Monte Carlo Markov Chain (MCMC) methods. Additionally, it is easier for users to adapt different forward models, and performs exceptionally well in convex landscapes. When landscapes are non-convex, however, the inverse Hessian may not provide a comprehensive view of the parameter space’s uncertainty; instead, uncertainty estimation could be computed using the variance across multiple BICePs runs starting from different initial parameters. The variational minimization approach also requires careful consideration of disperse starting parameters to ensure global minimization.
@@
${{}^{3}\!J}_{C^{\prime}C}$ 1 $0^{\circ}$ $1.71\pm 0.02$ $-0.85\pm 0.01$ $0.54\pm 0.00$
2 $0^{\circ}$ $1.30\pm 0.03$ $-0.91\pm 0.01$ $0.62\pm 0.01$
3 $0^{\circ}$ $1.62\pm 0.03$ $-0.87\pm 0.01$ $0.63\pm 0.01$
@@
${{}^{3}\!J}_{H^{\alpha}C^{\prime}}$ 1 $120^{\circ}$ $3.64\pm 0.02$ $-2.14\pm 0.02$ $1.27\pm 0.02$
2 $120^{\circ}$ $4.10\pm 0.03$ $-2.00\pm 0.02$ $0.95\pm 0.02$
3 $120^{\circ}$ $3.78\pm 0.02$ $-2.12\pm 0.02$ $1.21\pm 0.02$
${{}^{3}\!J}_{H^{N}C^{\prime}}$ 1 $180^{\circ}$ $4.33\pm 0.04$ $-1.17\pm 0.01$ $0.14\pm 0.01$
2 $180^{\circ}$ $4.60\pm 0.12$ $-0.57\pm 0.03$ $-0.10\pm 0.01$
3 $180^{\circ}$ $4.57\pm 0.09$ $-1.20\pm 0.03$ $0.13\pm 0.01$
${{}^{3}\!J}_{H^{N}C^{\beta}}$ 1 $60^{\circ}$ $2.72\pm 0.03$ $-0.35\pm 0.03$ $0.12\pm 0.01$
2 $60^{\circ}$ $3.00\pm 0.04$ $-0.26\pm 0.03$ $-0.28\pm 0.02$
3 $60^{\circ}$ $2.52\pm 0.03$ $-0.03\pm 0.02$ $-0.09\pm 0.02$
${{}^{3}\!J}_{H^{N}H^{\alpha}}$ 1 $-60^{\circ}$ $7.11\pm 0.05$ $-1.38\pm 0.03$ $1.43\pm 0.04$
2 $-60^{\circ}$ $7.50\pm 0.07$ $-1.50\pm 0.02$ $1.50\pm 0.06$
3 $-60^{\circ}$ $6.97\pm 0.07$ $-1.49\pm 0.04$ $1.63\pm 0.05$
@@
The BICePs score, $f_{\xi=0\rightarrow 1}^{\text{1d3z}}=38.14\pm 0.08$ (Table 2) is equivalent (within error) with the most probable landscape basin $f=38.15\pm 0.19$ from sampling the energy landscape, computed as an averarge across four chains; an example for one chain is shown in Figure 4. This is additional evidence of the algorithm’s reliability ,and corroborates that the results from variational minimization of the BICePs score and full joint posterior sampling are equivalent.
@@
Parameters $f_{\xi=0\rightarrow 1}^{\text{1d3z}}$ $f_{\xi=0\rightarrow 1}^{\text{2nr2}}$ $*f_{\xi=0\rightarrow 1}^{\text{CHARMM22*}}$
@@
Figure 6: A visualization of the 2NR2 conformational ensemble after joint conformational populations and Karplus parameter optimization using BICePs. Structures (144 in total) with higher transparency have lower populations. Residues are colored by the magnitude of the replica-averaged data deviation from the experiment in ${{}^{3}\!J}_{C^{\prime}C^{\prime}}$, where the color is normalized from blue (lowest deviation) to red (highest deviation).
It is difficult to say which of the model parameters are the best for ubiquitin, so we compare the top four: BICePs(RF2), Bax 1997, BICePs(1D3Z), and BICePs(2NR2). The BICePs(2NR2) parameters are objectively better at predicting $J$-couplings from structures of ubiquitin generated from simulations using CHARMM22* force field. Futhermore, our BICePs(RF2) parameters have slightly better transferability across structural ensembles and have a better BICePs score over Bax 1997 parameters. In general, when looking across structural ensembles, the lowest BICePs scores come from the 1D3Z structural ensemble ($f_{\xi=0\rightarrow 1}^{\text{1d3z}}$) except for BICePs parameters derived from the 2NR2 ensemble (BICePs(2NR2)). This confirms that the 1D3Z structural ensemble gives the strongest agreement with experimental NMR observations.

### v3

@@ Abstract
To quantify how well theoretical predictions of structural ensembles agree with experimental measurements, we depend on the accuracy of forward models. These models are computational frameworks that generate observable quantities from molecular configurations based on empirical relationships linking specific molecular properties to experimental measurements. Bayesian Inference of Conformational Populations (BICePs) is a reweighting algorithm that reconciles simulated ensembles with ensemble-averaged experimental observations, even when such observations are sparse and/or noisy. This is achieved by sampling the posterior distribution of conformational populations under experimental restraints as well as sampling the posterior distribution of uncertainties due to random and systematic error. In this study, we enhance the algorithm for the refinement of empirical forward model (FM) parameters. We introduce and evaluate two novel methods for optimizing FM parameters. The first method treats FM parameters as nuisance parameters, integrating over them in the full posterior distribution. The second method employs variational minimization of a quantity called the BICePs score that reports the free energy of “turning on” the experimental restraints. This technique, coupled with improved likelihood functions for handling experimental outliers, facilitates force field validation and optimization, as illustrated in recent studies (Raddi et al. 2023, 2024). Using this approach, we refine parameters that modulate the Karplus relation, crucial for accurate predictions of $J$-coupling constants based on dihedral angles ($\phi$) between interacting nuclei. We validate this approach first with a toy model system, and then for human ubiquitin, predicting six sets of Karplus parameters for ${{}^{3}\!J}_{H^{N}H^{\alpha}}$, ${{}^{3}\!J}_{H^{\alpha}C^{\prime}}$, ${{}^{3}\!J}_{H^{N}C^{\beta}}$, ${{}^{3}\!J}_{H^{N}C^{\prime}}$, ${{}^{3}\!J}_{C^{\prime}C^{\beta}}$, ${{}^{3}\!J}_{C^{\prime}C^{\prime}}$. Finally, we demonstrate that our framework naturally generalizes optimization to any differentiable forward model, such as those constructed with neural networks. This approach provides a promising direction for training and validating neural network-based forward models.
@@ I Introduction
Model validation and refinement of structural ensembles against NMR observables critically depends on reliable forward models (FMs) that have been robustly parameterized to reduce error. In refining these FMs, it is important to consider random and systematic errors inherent to the experimental data, which in turn contribute to overall uncertainty of integrative models. It is usually challenging to obtain complete knowledge of such errors; instead, such errors typically need to be inferred from the data.
@@
Beyond structural biology, Bayesian methods are widely used in survival analysis to model censored and heterogeneous discrete data. Recent work by Akhtar and colleagues shows that flexible priors and robust loss functions enable accurate parameter estimation and outlier resistance in such settings Akhtar et al. (2024, 2025). These developments highlight the versatility of Bayesian inference for complex, noisy datasets and provide additional examples of analogous strategies for handling heterogeneity and outliers in our domain.
To effectively refine FM parameters, we use BICePs to sample over the entire posterior distribution of FM parameters, enabling simultaneous ensemble reweighting and parameter refinement. Additionally, we show that by variational minimization of the BICePs score, we obtain the same result and show that the two approaches are equivalent, with each method requiring particular considerations. We first demonstrate our method’s effectiveness using a toy model and then apply it to optimize six distinct sets of Karplus parameters for the human protein ubiquitin. As a proof-of-concept for more complex models, we also train a neural network to learn $J$-coupling constants from dihedral angles using the BICePs score as a loss function. We then discuss how this idea could be applied to neural networks for chemical shift prediction.Han et al. (2024); Yang, Chakraborty, and White (2021); Meiler (2003); Shen and Bax (2010) Of course, our framework naturally extends to other empirical forward models with tunable parameters—for example, hydrogen-deuterium exchange (HDX) models with adjustable hydrogen-bonding and burial weights, or paramagentic relaxation enhancement (PRE) models with system-specific correlation times and scaling factors. These examples illustrate the generality and flexibility of our approach for refining a wide range of forward models to improve the agreement between simulations and experiment.
@@ II Theory
In this section, we develop two different theoretical approaches to optimize forward models. In the first method, we include forward model parameters in the posterior, and sample the full posterior distribution of all parameters. In the second approach, we optimize the forward model by minimizing the BICePs score, a free energy-like quantity that can be used for variational optimization. While the two methods are theoretically equivalent (see the supplemental material for a short proof of this), each has practical advantages and disadvantages.
@@ II.1 Forward model optimization by posterior sampling of forward model parameters
The Good-and-Bad model accounts for systematic error due to outlier measurements.
Recent work has developed sophisticated likelihood models for BICePs, designed to detect outliers, that are robust in the presence of systematic errorRaddi et al. (2025). Raddi et al. (2025) developed two likelihood models for this purpose: the Student’s model and the Good-and-Bad model. In related work using BICePs to perform automated force field optimization, we used the Student’s model.Raddi and Voelz (2024) Here, in this work, we use the Good-and-Bad model, primarily because the derivatives are less complicated than the Student’s model.
The Good-and-Bad likelihood modelRaddi et al. (2025) assumes that the level of noise is mostly uniform, except for a few erratic measurements. This limits the number of uncertainty parameters that need to be sampled, while still capturing outliers. Consider a model in which the uncertainties $\sigma_{j}$ for particular observables $j$ are distributed with respect to some typical uncertainty $\sigma^{B}$ according to a conditional probability $p(\sigma_{j}|\sigma^{B})$. We derive a posterior for the parameters of the $k^{th}$ FM having a single uncertainty parameter $\sigma^{B}$ by marginalizing over all $\sigma_{j}$,
$p(\mathbf{X},\sigma_{0},\theta^{(k)}|D)\propto\prod^{N}_{r=1}p(X_{r})\prod^{N_{d}}_{j=1}\int\limits_{\sigma^{SEM}}^{\infty}p(d_{j}|g_{j}(\mathbf{X},\theta^{(k)}),\sigma_{j})p(\sigma_{j}|\sigma_{0})d\sigma_{j}$ (7)
@@
where $0\leq\omega<1$ describes the fraction of "bad" observables. Since the value of $\omega$ is unknown, it is treated as a nuisance parameter, and marginalized over its range. The resulting posterior is
$\begin{split}&p(\mathbf{X},\sigma_{0},\varphi,\theta^{(k)}|D)\\ &\propto\prod^{N}_{r=1}\left\{p(X_{r})\prod^{N_{d}}_{j=1}\int\limits_{0}^{1}d\omega\int\limits_{\sigma^{\text{SEM}}}^{\infty}\exp\left(-\frac{(d_{j}-g_{j}(\mathbf{X},\theta^{(k)}))^{2}}{2\sigma_{j}^{2}}\right)\frac{\omega\delta(\sigma_{j}-\varphi\sigma_{0})+(1-\omega)\delta(\sigma_{j}-\sigma_{0})}{\sqrt{2\pi}\sigma_{j}}d\sigma_{j}\right\}\\ &=\prod^{N}_{r=1}\left\{p(X_{r})\prod^{N_{d}}_{j=1}\Biggl(\frac{1-H(\sigma^{SEM}-\sigma_{0})}{2}\mathcal{N}\left(g_{j}(\mathbf{X},\theta^{(k)})\,\middle|\,d_{j},\sigma_{0}^{2}\right)+\frac{1-H(\sigma^{SEM}-\varphi\sigma_{0})}{2\varphi}\mathcal{N}\left(g_{j}(\mathbf{X},\theta^{(k)})\,\middle|\,d_{j},\varphi^{2}\sigma_{0}^{2}\right)\Biggr)\right\},\end{split}$ (9)
@@
$\begin{split}\hskip-7.11317ptu=\sum^{N}_{r=1}-\log(p(X_{r}))-N\sum^{N_{d}}_{j=1}\log\left[\frac{1-H(\sigma^{SEM}-\sigma_{0})}{2}\mathcal{N}\left(g_{j}(\mathbf{X},\theta^{(k)})\,\middle|\,d_{j},\sigma_{0}^{2}\right)+\frac{1-H(\sigma^{SEM}-\varphi\sigma_{0})}{2\varphi}\mathcal{N}\left(g_{j}(\mathbf{X},\theta^{(k)})\,\middle|\,d_{j},\varphi^{2}\sigma_{0}^{2}\right)\right].\end{split}$ (10)
@@ II.2 Forward model optimization by variational minimization of the BICePs score
As we show in Results and Discussion, treating the forward model parameters as nuisance parameters and sampling them as part of the full posterior is an efficient strategy to refine FM parameters. However, in the limit of a very large number of FM parameters, the dimensionality of the posterior may become unwieldy and may suffer from the “curse of dimensionality”.
@@
$Z(\theta)=\iint\exp\left(-u(\mathbf{X},\mathbf{\sigma};\theta)\right)d\mathbf{X}d\mathbf{\sigma}$ (16)
@@
It should be noted that in other applications of BICePs,Raddi et al. (2025); Raddi and Voelz (2024) the reference state for the BICePs score is defined using the $\lambda=0$ state for a series of priors $p_{\lambda}(X)\sim[p(X)]^{\lambda}$, and the BICePs score is computed as the free energy of ($\lambda=0\rightarrow 1$) and ($\xi=0\rightarrow 1$) transformations. Here, since we are only interested in evaluating and/or parameterizing the likelihood functions, we set $p(X)$ to be uniform.
@@
For this purpose, gradient-based searches are very helpful. Conveniently, the derivative of the BICePs score $f(\theta)$ with respect to the FM parameters reduces to the Boltzmann-averaged values of the gradient $\partial u/\partial\theta$ :
$\begin{split}\frac{\partial f(\theta)}{\partial\theta_{i}}&=\iint\frac{1}{Z(\theta)}\left[\frac{\partial u}{\partial\theta_{i}}\right]\exp\left(-u\right)d\mathbf{X}d\mathbf{\sigma}=\bigg\langle\frac{\partial u}{\partial\theta_{i}}\bigg\rangle\end{split}$ (17)
@@ III Results
To evaluate the robustness of our algorithm in the presence of experimental inaccuracies, we introduced random and systematic errors of varying magnitudes ($\sigma_{\text{data}}$) into the synthetic experimental scalar couplings. The performance of our Good-Bad likelihood model, a Gaussian likelihood model, and singular value decomposition (SVD) was compared under these conditions.
@@
Parameters of many reference Karplus relations were previously determined using a singular value determination approach.Hu and Bax (1997) To test this approach against our BICePs method, we derived Karplus parameters $\theta=\{$A,B,C}\} using a weighted singular value decomposition (SVD) fitting approach to optimally fit the $J$-coupling values as a function of dihedral angles. For each observation $j$ across $N_{d}$ measurements, the matrix $M$ was constructed with rows for each $\phi$ angle:
@@
where $p(X)$ represents the true populations for state $X$, and $\phi_{0}$ is the phase shift of $60^{\circ}$.
@@
Figure 2: Comparative analysis in the performance of the Good-Bad likelihood model (red), a Gaussian likelihood model (blue), and singular value decomposition (SVD) using the "true" $\phi$ angles with synthetic experimental data. Here, we induced random and systematic error of varying magnitude ($\sigma_{\text{data}}$) to the experimental scalar couplings. Model performance was measured by computing RMSE (Hz) between the "true" scalar couplings and the couplings generated from the Karplus relations with predicted Karplus coefficients over 1,500 random perturbations to the experimental data, and represent the average of 100 BICePs calculations. Error bars represent the standard deviation. Predictions from SVD and the Gaussian likelihood model become notably less dependable when data incorporates errors, especially when $\sigma_{\text{data}}$ exceeds 0.5 Hz.
@@
As a method for forward model optimization, variational minimization of the BICePs score has advantages and disadvantages. This method is particularly advantageous for handling many FM parameters, offering a potential solution to the curse of dimensionality faced by basic Monte Carlo Markov Chain (MCMC) methods. Additionally, it is easier for users to adapt different forward models, and performs exceptionally well in convex landscapes. When landscapes are non-convex, however, the inverse Hessian may not provide a comprehensive view of the parameter space’s uncertainty; instead, uncertainty estimation could be computed using the variance across multiple BICePs runs starting from different initial parameters. The variational minimization approach also requires careful consideration of disperse starting parameters to ensure global minimization.
@@
${{}^{3}\!J}_{C^{\prime}C}$ 1 $0^{\circ}$ $1.71\pm 0.02$ $0.85\pm 0.01$ $0.54\pm 0.00$
2 $0^{\circ}$ $1.30\pm 0.03$ $0.91\pm 0.01$ $0.62\pm 0.01$
3 $0^{\circ}$ $1.62\pm 0.03$ $0.87\pm 0.01$ $0.63\pm 0.01$
@@
${{}^{3}\!J}_{H^{\alpha}C^{\prime}}$ 1 $120^{\circ}$ $3.64\pm 0.02$ $2.14\pm 0.02$ $1.27\pm 0.02$
2 $120^{\circ}$ $4.10\pm 0.03$ $2.00\pm 0.02$ $0.95\pm 0.02$
3 $120^{\circ}$ $3.78\pm 0.02$ $2.12\pm 0.02$ $1.21\pm 0.02$
${{}^{3}\!J}_{H^{N}C^{\prime}}$ 1 $180^{\circ}$ $4.33\pm 0.04$ $1.17\pm 0.01$ $0.14\pm 0.01$
2 $180^{\circ}$ $4.60\pm 0.12$ $0.57\pm 0.03$ $0.10\pm 0.01$
3 $180^{\circ}$ $4.57\pm 0.09$ $1.20\pm 0.03$ $0.13\pm 0.01$
${{}^{3}\!J}_{H^{N}C^{\beta}}$ 1 $60^{\circ}$ $2.72\pm 0.03$ $0.35\pm 0.03$ $0.12\pm 0.01$
2 $60^{\circ}$ $3.00\pm 0.04$ $0.26\pm 0.03$ $0.28\pm 0.02$
3 $60^{\circ}$ $2.52\pm 0.03$ $03\pm 0.02 0.0$$09\pm 0.02$
${{}^{3}\!J}_{H^{N}H^{\alpha}}$ 1 $60^{\circ}$ $7.11\pm 0.05$ $1.38\pm 0.03$ $1.43\pm 0.04$
2 $60^{\circ}$ $7.50\pm 0.07$ $1.50\pm 0.02$ $1.50\pm 0.06$
3 $60^{\circ}$ $6.97\pm 0.07$ $1.49\pm 0.04$ $1.63\pm 0.05$
@@
The BICePs score, $f_{\xi=0\rightarrow 1}^{\text{1d3z}}=38.14\pm 0.08$ (Table 2) is equivalent (within error) with the most probable landscape basin $f=38.15\pm 0.19$ from sampling the energy landscape, computed as an average across four chains; an example for one chain is shown in Figure 4. This is additional evidence of the algorithm’s reliability, and corroborates that the results from variational minimization of the BICePs score and full joint posterior sampling are equivalent.
@@
Parameters $f_{\xi=0\rightarrow 1}^{\text{1d3z}}$ $f_{\xi=0\rightarrow 1}^{\text{2nr2}}$ $f_{\xi=0\rightarrow 1}^{\text{CHARMM22*}}$
@@
We obtained Karplus parameters using SVD with uniform populations across the 1D3Z ensemble (Table S1), denoted SVD(1D3Z). As shown in Table 2, these parameters perform well for 1D3Z itself ($f_{\xi=0\rightarrow 1}^{\text{1d3z}}=44.76\pm 0.66$) but generalize poorly to other ensembles, giving higher scores for 2NR2 and CHARMM22*. The MAE results in Table S2 show a similar trend: while SVD(1D3Z) predictions are competitive within 1D3Z, they are consistently worse than BICePs(1D3Z) for 2NR2 and CHARMM22*. This highlights a key limitation of SVD with uniform populations—without proper treatment of ensemble populations, the parameters risk overfitting to the training ensemble and lose transferability. By contrast, BICePs jointly optimizes both parameters and populations, yielding lower errors across ensembles and more robust performance.
Figure 6: A visualization of the 2NR2 conformational ensemble after joint conformational populations and Karplus parameter optimization using BICePs. Structures (144 in total) with higher transparency have lower populations. Residues are colored by the magnitude of the deviation between the replica-averaged forward model observations $g(\mathbf{X},\theta)$ and the experimental ${{}^{3}\!J}_{C^{\prime}C^{\prime}}$ data $d$, where the color is normalized from blue (lowest deviation) to red (highest deviation).
It is difficult to say which of the model parameters are the best for ubiquitin, so we compare the top four: BICePs(RF2), Bax 1997, BICePs(1D3Z), and BICePs(2NR2). The BICePs(2NR2) parameters are objectively better at predicting $J$-couplings from structures of ubiquitin generated from simulations using CHARMM22* force field. In general, when looking across structural ensembles, the lowest BICePs scores come from the 1D3Z structural ensemble ($f_{\xi=0\rightarrow 1}^{\text{1d3z}}$) except for BICePs parameters derived from the 2NR2 ensemble (BICePs(2NR2)). This confirms that the 1D3Z structural ensemble gives the strongest agreement with experimental NMR observations. However, our BICePs(RF2) parameters have the best transferability across structural ensembles and have lower BICePs scores compared to Bax1997 parameters. In a separate study, we evaluated the performance of the BICePs(RF2) parameters (denoted as “Raddi2024” in refNguyen, Raddi, and Voelz (2025)) in predicting ${{}^{3}\!J}_{H^{N}H^{\alpha}}$ couplings for a set of non-natural and cyclic $\beta$-hairpin peptidesNguyen, Raddi, and Voelz (2025). Our results show that these parameters yield consistently lower prediction errors compared to the Bax2007Vögeli et al. (2007) and Kessler1988Kessler et al. (1988) parameter sets. Based on all evidence, we recommend the BICePs(RF2) parameters for downstream applications such as structural refinement and validation. Nonetheless, users should carefully assess accuracy for their specific system, as additional refinement may still be required.